Broussard Skateboard's sales are expected to increase by 25% from $8.4 million in 2016 to $10.50 million in 2017. Its assets totaled $3 million at the end of 2016. Broussard is already at full capacity, so its assets must grow at the same rate as projected sales. At the end of 2016, current liabilities were $1.4 million, consisting of $450,000 of accounts payable, $500,000 of notes payable, and $450,000 of accruals. The after-tax profit margin is forecasted to be 4%. Assume that the company pays no dividends. Under these assumptions, what would be the additional funds needed for the coming year? Do not round intermediate calculations. Round your answer to the nearest dollar.
Answer #1

AFN = Increase in total assets - Increase in liabilities - Addition to retained earnings

Increase in total assets = Assets * Growth rate = 3000000 * 25% = 750000

Spontaneous liabilities = Accounts payable + accruals = 450000 + 450000 = 900000

Increase in spontaneous liabilities = 900000 * growth rate = 900000 * 25% = 225000

Increase in retained earnings = Profit margin * retention ratio * forecasted sales

= 4% * (100% -0%) * 10500000

= 420000

AFN = 750000 - 225000 - 420000

= 105000